OpenStudy (anonymous):

Ace Electronics, Hi-Fi Stereo, and CD World buy a total of 170 VCR/DVD players from the same distributor each month. Because CD World buys 30 more units than the other two stores combined, CDW's cost is only $160 per unit. The players cost Hi-Fi $165 each and Ace $170 each. How many players does each retailer buy each month if the distributor receives $27,675 each month from the sale of the players to the three stores?

OpenStudy (anonymous):

sorry not good at math

OpenStudy (anonymous):

AE 170 per unit HiFi 165 per unit CDW 30 more units than other 2 @ 160 per unit 27675 = 170a + 165h + 160c(30) I THINK you just now go thru & solve for each letter. Im not a math whiz, but this is what I would do if I had this problem.

OpenStudy (anonymous):

AE = # of discs Ace Electronics buys HF = # of discs Hi-Fi Stereo buys CD = # of discs CD World buys (1) AE + HF + CD = 170 (2) CD = (AE + HF) + 30 (3) 160CD + 165HF + 170AE = 27675 3 Equations. 3 Unknowns. Solve for each.